How many grams of iron is in 6.42 × 10E-2 mol Fe?
Answer #1

Mass of 1 mole of Fe = 55.845 g

mass of 6.42*10^-2 mole of Fe =

= 6.42*10^-2 mole of Fe *(55.845 g / 1 mole of Fe)

= 359 * 10^-2 g

= 3.59 g

Therefore, the mass of 6.42*10^-2 mole of Fe = 3.59 g